OpenStudy (anonymous):

The probability of Keith winning a chess game against Sooeng is 0.4. What is the probability that Keith will not win 3 games in a row?

OpenStudy (anonymous):

\[1-\left(\frac{4}{10}\right)^3=\frac{117}{125}=0.936 \]

OpenStudy (anonymous):

so all the probabilities (100%) - the chance of 3 wins in a row (win chance)*(win chance)*(win chance) = 1-(0.4)^3